OpenStudy (anonymous):

proof that cos 2x= 2 ((cos)^2(x))-1

OpenStudy (anonymous):

cos 2x = cos (x + x) by the sum rule, cos (x + x) = cosx*cosx - sinx*sinx = cos^2(x) - sin^2(x) since sin^2(x) = 1-cos^2(x) cos^2(x) - sin^2(x) = cos^2(x) - (1 - cos^2(x)) = cos^2(x) - 1 + cos^2(x)) = 2*cos^2(x) - 1

OpenStudy (anonymous):

i got it, thank you